OpenStudy (anonymous):

Find the real-number solutions of the equation? 27g^3-8=0

OpenStudy (anonymous):

27g^3-8 = 0 27g^3 = 8 g^3 = 8/27 g = (8/27)^(1/3) cube root of 8 is 2 cube root of 27 is 3 g = 2/3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcome